OpenStudy (anonymous):

which of the following equals 7^2 A.10 B.49 C.50 D.75

OpenStudy (anonymous):

b.

OpenStudy (ksaimouli):

b 49

OpenStudy (ksaimouli):



OpenStudy (anonymous):

b

OpenStudy (anonymous):

7*7=49

OpenStudy (ksaimouli):

7*1=7 7*2=14 ....... 7*7=49

OpenStudy (ksaimouli):

did u get it

OpenStudy (anonymous):

i get it i just mixed up with the signs and numbers